OpenStudy (anonymous):

For an angle Θ with the point (12, −5) on its terminating side, what is the value of cosine?

OpenStudy (anonymous):



OpenStudy (anonymous):

ok then what? sorry I just taught myself literally 15 minutes ago

OpenStudy (anonymous):

I know how to set it up but I dontknow what to do next

OpenStudy (anonymous):

\[Cos(\theta) =\frac{ Adjacent }{ Hypotenuse }\] The terminal side is the hypotenuse. \[Hypotenuse = \sqrt{A^2+B^2}\] \[Hypotenuse = \sqrt{(12)^2+(-5)^2}\] Solve for the Hypotenuse and then use the Adjacent side, the side adjacent to your angle, which is 12 in your case.

OpenStudy (anonymous):

You should be able to solve from here.

OpenStudy (anonymous):

so √144+25 √169 13? but that's not an option

OpenStudy (anonymous):

That is only your hypotenuse. You must find the value of cosine by relating it to your formula for cos(angle)

OpenStudy (anonymous):

ohhh so then what?

OpenStudy (anonymous):

Relate it to your cosine(angle) formula. Adjacent side over Hypotenuse.

OpenStudy (anonymous):

how do I find adj?

OpenStudy (anonymous):

It is given.

OpenStudy (anonymous):

oh so 12/13?

OpenStudy (anonymous):